Lee Company manufactures a product for which the following direct material

standards have been set: 3 feet per unit at a standard unit cost of $15. During

October, the company purchased direct materials at a cost of $55,650, all of which

were used in the production of 3,200 units of product. In addition, 4,900 hours of

direct labor were logged during the month. The cost of labor time was $36,750. For

the month, the direct materials quantity variance was an unfavorable $4,500. The

total labor variances were a favorable $1,650; the direct labor efficiency variance

was an unfavorable $800.

Calculate the following:

1.

The actual cost per foot of material for October
Answer #1
Standard cost per foot of materials = Standard unit materials cost / Standard direct materials per unit = 15 / 3 5
Standard quantity of materials for actual production = Actual units produced * Standard quantity of direct materials per unit = 3200 * 3 9600
Unfavorable variance is written as positive value
Direct materials quantity variance = ( Actual quantity - Standard quantity ) * Standard price
4500 = ( Actual quantity - 9600 ) * 5
4500 / 5 = Actual quantity - 9600
900 = Actual quantity - 9600
Actual quantity = 900 + 9600 10500
Actual cost per foot of materials = Actual cost of materials / Actual quantity of materials = 55650 / 10500    5.3